### v3

@@ 4.3 Which ranking method is most effective?
FlanT5-large TI_3, OT_1, TW_0, PF, B, RP_1 TI_1, OT_1, TW_4, QF, B, RP_None TI_3, OT_2, TW_2, PF, E, RP_None TI_1, OT_3, TW_0, QF, B, RP_None
FlanT5-xl TI_2, OT_3, TW_4, PF, E, RP_1 TI_1, OT_1, TW_2, QF, B, RP_1 TI_3, OT_2, TW_3, QF, B, RP_None TI_1, OT_3, TW_3, QF, E, RP_1
FlanT5-xxl TI_4, OT_3, TW_1, PF, B, RP_None TI_1, OT_1, TW_0, QF, E, RP_1 TI_1, OT_2, TW_2, QF, B, RP_None TI_1, OT_1, TW_4, QF, B, RP_1
Mistral-7B TI_1, OT_4, TW_3, PF, E, RP_1 TI_1, OT_1, TW_0, QF, E, RP_1 TI_1, OT_1, TW_2, QF, B, RP_None TI_1, OT_3, TW_0, QF, E, RP_1
Llama3-8B TI_2, OT_3, TW_2, PF, B, RP_1 TI_1, OT_1, TW_0, QF, B, RP_None TI_1, OT_2, TW_0, QF, B, RP_None TI_1, OT_1, TW_3, PF, E, RP_None
@@
FlanT5-large TI_3, OT_3, TW_3, PF, B, RP_1 TI_1, OT_1, TW_2, PF, E, RP_None TI_3, OT_2, TW_1, PF, E, RP_None TI_1, OT_1, TW_2, PF, E, RP_1
FlanT5-xl TI_3, OT_3, TW_4, QF, B, RP_1 TI_1, OT_1, TW_5, QF, B, RP_1 TI_2, OT_2, TW_3, PF, E, RP_1 TI_1, OT_2, TW_5, PF, B, RP_None
FlanT5-xxl TI_3, OT_1, TW_3, PF, B, RP_None TI_1, OT_1, TW_2, PF, E, RP_1 TI_2, OT_2, TW_3, PF, B, RP_None TI_1, OT_3, TW_3, PF, B, RP_1
Mistral-7B TI_1, OT_4, TW_2, PF, E, RP_1 TI_1, OT_1, TW_1, QF, E, RP_None TI_3, OT_1, TW_0, QF, B, RP_1 TI_1, OT_3, TW_2, QF, E, RP_1
Llama3-8B TI_1, OT_3, TW_1, PF, B, RP_1 TI_1, OT_1, TW_3, PF, E, RP_1 TI_3, OT_2, TW_0, QF, B, RP_None TI_1, OT_2, TW_4, PF, E, RP_None
@@
FlanT5-large TI_3, OT_1, TW_1, PF, B, RP_1 TI_1, OT_1, TW_5, PF, B, RP_1 TI_3, OT_2, TW_4, PF, E, RP_None TI_1, OT_1, TW_1, PF, B, RP_None
FlanT5-xl TI_3, OT_1, TW_1, PF, B, RP_1 TI_1, OT_1, TW_5, PF, B, RP_1 TI_2, OT_2, TW_1, QF, E, RP_None TI_1, OT_3, TW_5, QF, E, RP_None
FlanT5-xxl TI_3, OT_1, TW_1, PF, B, RP_None TI_1, OT_1, TW_0, PF, B, RP_1 TI_1, OT_2, TW_3, PF, B, RP_None TI_1, OT_3, TW_5, PF, B, RP_1
Mistral-7B TI_2, OT_2, TW_4, PF, B, RP_1 TI_1, OT_1, TW_3, QF, B, RP_1 TI_1, OT_1, TW_4, QF, B, RP_None TI_1, OT_1, TW_2, QF, B, RP_None
Llama3-8B TI_2, OT_3, TW_5, QF, E, RP_1 TI_1, OT_1, TW_4, QF, B, RP_None TI_1, OT_2, TW_0, QF, B, RP_None TI_1, OT_3, TW_1, QF, B, RP_1
@@ 6 Conclusions
6.0.1 Acknowledgment.
Shuoqi Sun is partially supported by the Australian Research Council through the ARC Centre of Excellence for Automated Decision-Making and Society (ADM+S, CE200100005).

### v4

@@ Abstract
Publication Notice: This paper has been accepted for publication at the 47th European Conference on Information Retrieval (ECIR 2025). The published version is available in the conference proceedings.
@@ 4.3 Which ranking method is most effective?
FlanT5-large
TI_3, OT_1,
TW_0, PF, B, RP_1
TI_1, OT_1,
TW_4, QF, B, RP_None
TI_3, OT_2,
TW_2, PF, E, RP_None
TI_1, OT_3,
TW_0, QF, B, RP_None
FlanT5-xl
TI_2, OT_3,
TW_4, PF, E, RP_1
TI_1, OT_1,
TW_2, QF, B, RP_1
TI_3, OT_2,
TW_3, QF, B, RP_None
TI_1, OT_3,
TW_3, QF, E, RP_1
FlanT5-xxl
TI_4, OT_3,
TW_1, PF, B, RP_None
TI_1, OT_1,
TW_0, QF, E, RP_1
TI_1, OT_2,
TW_2, QF, B, RP_None
TI_1, OT_1,
TW_4, QF, B, RP_1
Mistral-7B
TI_1, OT_4,
TW_3, PF, E, RP_1
TI_1, OT_1,
TW_0, QF, E, RP_1
TI_1, OT_1,
TW_2, QF, B, RP_None
TI_1, OT_3,
TW_0, QF, E, RP_1
Llama3-8B
TI_2, OT_3,
TW_2, PF, B, RP_1
TI_1, OT_1,
TW_0, QF, B, RP_None
TI_1, OT_2,
TW_0, QF, B, RP_None
TI_1, OT_1,
TW_3, PF, E, RP_None
@@
FlanT5-large
TI_3, OT_3,
TW_3, PF, B, RP_1
TI_1, OT_1,
TW_2, PF, E, RP_None
TI_3, OT_2,
TW_1, PF, E, RP_None
TI_1, OT_1,
TW_2, PF, E, RP_1
FlanT5-xl
TI_3, OT_3,
TW_4, QF, B, RP_1
TI_1, OT_1,
TW_5, QF, B, RP_1
TI_2, OT_2,
TW_3, PF, E, RP_1
TI_1, OT_2,
TW_5, PF, B, RP_None
FlanT5-xxl
TI_3, OT_1,
TW_3, PF, B, RP_None
TI_1, OT_1,
TW_2, PF, E, RP_1
TI_2, OT_2,
TW_3, PF, B, RP_None
TI_1, OT_3,
TW_3, PF, B, RP_1
Mistral-7B
TI_1, OT_4,
TW_2, PF, E, RP_1
TI_1, OT_1,
TW_1, QF, E, RP_None
TI_3, OT_1,
TW_0, QF, B, RP_1
TI_1, OT_3,
TW_2, QF, E, RP_1
Llama3-8B
TI_1, OT_3,
TW_1, PF, B, RP_1
TI_1, OT_1,
TW_3, PF, E, RP_1
TI_3, OT_2,
TW_0, QF, B, RP_None
TI_1, OT_2,
TW_4, PF, E, RP_None
@@
FlanT5-large
TI_3, OT_1,
TW_1, PF, B, RP_1
TI_1, OT_1,
TW_5, PF, B, RP_1
TI_3, OT_2,
TW_4, PF, E, RP_None
TI_1, OT_1,
TW_1, PF, B, RP_None
FlanT5-xl
TI_3, OT_1,
TW_1, PF, B, RP_1
TI_1, OT_1,
TW_5, PF, B, RP_1
TI_2, OT_2,
TW_1, QF, E, RP_None
TI_1, OT_3,
TW_5, QF, E, RP_None
FlanT5-xxl
TI_3, OT_1,
TW_1, PF, B, RP_None
TI_1, OT_1,
TW_0, PF, B, RP_1
TI_1, OT_2,
TW_3, PF, B, RP_None
TI_1, OT_3,
TW_5, PF, B, RP_1
Mistral-7B
TI_2, OT_2,
TW_4, PF, B, RP_1
TI_1, OT_1,
TW_3, QF, B, RP_1
TI_1, OT_1,
TW_4, QF, B, RP_None
TI_1, OT_1,
TW_2, QF, B, RP_None
Llama3-8B
TI_2, OT_3,
TW_5, QF, E, RP_1
TI_1, OT_1,
TW_4, QF, B, RP_None
TI_1, OT_2,
TW_0, QF, B, RP_None
TI_1, OT_3,
TW_1, QF, B, RP_1